# Narrating the Video: Boosting Text-Video Retrieval via Comprehensive Utilization of Frame-Level Captions

### v1

@@ B.2 Implementation Details
C Caption Generator
When using large multimodal models (LMMs) for caption generation, both the caption quality and inference time can vary depending on the prompt. Although our NarVid framework generates narrations offline, the computational resources and time required for pre-generating captions remain practical challenges in our experiments. Finding the optimal prompt for each dataset requires substantial experimentation, which may be beyond the scope of our current research. To address this challenge, we utilized MSVD dataset [3] as a testbed, which requires the least computational resources for inference among the datasets. We first evaluated the retrieval performance across various prompts on MSVD and then applied our findings to other datasets.
@@ C Caption Generator
As shown in Tab. 7, we provide several examples of prompts with their specific characteristics. We used the LLaVa 1.5 7B model [21] as a baseline model for the experiments. The prompts used for caption generation and their specific characteristics are shown in Tab. 7. Prompt A is a simple request to describe the objects within each frame in a single sentence. Prompt B is more structured, requiring a single-sentence description that includes objects, actions, and other parts. Prompt C removes the single-sentence constraint, allowing for a more detailed description of each frame. As shown in Tab. 8, prompt C achieves the highest performance in R@1, and we use it for other experiments. Given the apparent potential for improvement in these captions, we expect that various advanced inference techniques, such as in-context learning, will enhance the performance in future work. The captions used for experiments will be publicly available with our code.
@@ D Ablation Study
Table 9: Results depend on similarity matrices on MSR-VTT for zero-shot retrieval evaluation. Setting: we extracted CLIP features from the videos and their narrations to compare with the queries’ CLIP features without any training. $\boldsymbol{S}_{qv}$ and $\boldsymbol{S}_{qn}$ are the query-video and query-narration similarity matrices, each. $\boldsymbol{S}_{sum}$ indicates the element-wise sum of the two matrices, while $\boldsymbol{S}_{fusion}$ represents the element-wise sum of the two normalized matrices.
@@ D.1 Effectiveness of narration utilization
Zero-shot retrieval performance. Can captions from good models such as LMMs deliver outstanding results on their own? To answer this question, in Tab. 9, we evaluate zero-shot retrieval to verify the importance of effectively utilizing narration information without training. For the experiment setting, we extract CLIP features from each video and its narration, applied mean pooling, and evaluated zero-shot retrieval performance with given queries. In the table, the query-to-video retrieval shows reasonable retrieval performance without training, while the query-to-narration retrieval shows significantly lower performance. Although an element-wise matrix summation or the fusion approach used in our NarVid framework shows some improvements, the retrieval performances remain insufficient for practical retrieval scenarios. These experimental results once again support the claim of our framework that the information from the generative model should be utilized appropriately, including training.
@@
D.2 Cross-View Hard Negative Loss
Figure 7: Ablation study results of $\lambda$ and $\eta$, used in the hard negative loss, with $\alpha$ set to 1, conducted on the MSR-VTT
Fig. 7 shows the performance of R@1 with varying values of $\lambda$ and $\eta$, which are hyperparameters used in the hard negative rank loss of Eqs. 6 and 8. $\lambda$ is a scaling factor that establishes the threshold for selecting hard negatives within the query-video and query-narration similarity matrices. Likewise, $\eta$ serves as a scaling factor that modifies the thresholds in the cross-view hard negative loss. The balance between $\lambda$ and $\eta$ is crucial for optimal performance, achieving its best results at $\lambda=0.7$ and $\eta=1.8$.
@@ D.2 Cross-View Hard Negative Loss
$L_{qv}^{H}=-\frac{1}{2B}\sum_{i=1}^{B}\left\{\log\frac{e^{\boldsymbol{S}_{qv}(i,i)}}{\displaystyle\sum_{j=1}^{H_{i}}e^{\boldsymbol{S}_{qv}(i,j)}}+\log\frac{e^{\boldsymbol{S}_{qv}(i,i)}}{\displaystyle\sum_{j=1}^{H_{i}^{T}}e^{\boldsymbol{S}_{qv}(j,i)}}\right\},$
$L_{qn}^{H}=-\frac{1}{2B}\sum_{i=1}^{B}\left\{\log\frac{e^{\boldsymbol{S}_{qn}(i,i)}}{\displaystyle\sum_{j=1}^{H_{i}}e^{\boldsymbol{S}_{qn}(i,j)}}+\log\frac{e^{\boldsymbol{S}_{qn}(i,i)}}{\displaystyle\sum_{j=1}^{H_{i}^{T}}e^{\boldsymbol{S}_{qn}(j,i)}}\right\},$

### v4

@@ B.2 Implementation Details
Figure 6: Narration comparison on MSR-VTT across different LLaVA versions.
C Caption Generator
C.1 Prompt Strategies
When using large multimodal models (LMMs) for caption generation, both the caption quality and inference time can vary depending on the prompt. Although our NarVid framework generates narrations offline, the computational resources and time required for pre-generating captions remain practical challenges in our experiments. Finding the optimal prompt for each dataset requires substantial experimentation, which may be beyond the scope of our current research. To address this challenge, we utilized the MSVD dataset [3] as a testbed, which requires the least computational resources for inference among the datasets. We first evaluated the retrieval performance across various prompts on MSVD and then applied our findings to other datasets.
Captioner Time(s) Text-to-video retrieval
R@1 R@5 R@10
LLaVA 0.5B 7.92 50.0 76.2 84.9
LLaVA 7B 25.31 51.0 76.4 85.2
Table 9: Comparison of text-to-video retrieval results on MSR-VTT based on the LLaVA model size.
As shown in Tab. 7, we provide several examples of prompts with their specific characteristics. We used the LLaVa 1.5 7B model [22] as a baseline model for the experiments. Prompt A is a simple request to describe the objects within each frame in a single sentence. Prompt B is more structured, requiring a single-sentence description that includes objects, actions, and other parts. Prompt C removes the single-sentence constraint, allowing for a more detailed description of each frame. As shown in Tab. 8, prompt C achieves the highest performance in R@1, and we use it for other experiments. Given the apparent potential for improvement in these captions, we expect that various advanced inference techniques, such as in-context learning, will enhance the performance in future work. The captions used for experiments will be publicly available with our code.
C.2 Time Efficiency for Caption Generation
In our experiments, we generated all frame-level captions per video offline. As shown in Tab. 9, we mainly use the basic LLaVa 1.5 7B model as a captions generator, therefore the inference time can be reduced with inference boosting toolkits. One interesting point is that a lighter frame captioner (0.5B) can achieve considerable performance with less computing time.
C.3 Performance Discrepancy by LLaVA 1.5 7B vs 1.6 13B
In general, the larger VLM tends to generate more detail and long captions, but all of them are not necessarily related to given queries, which shows less portion of query vocabulary (Fig. 6 (a)). The well-generated captions with query-irrelevant words may cause a negative effect on the retrieval performance, as shown in Fig. 6 (b).
@@ D Ablation Study
Table 10: Results depending on similarity matrices on MSR-VTT for zero-shot retrieval evaluation. Setting: we extracted CLIP features from the videos and their narrations to compare with the queries’ CLIP features without any training. $\boldsymbol{S}_{qv}$ and $\boldsymbol{S}_{qn}$ are the query-video and query-narration similarity matrices, each. $\boldsymbol{S}_{sum}$ indicates the element-wise sum of the two matrices, while $\boldsymbol{S}_{fusion}$ represents the element-wise sum of the two normalized matrices.
@@ D.1 Effectiveness of Narration Utilization
Zero-shot retrieval performance. Can captions from good models such as LMMs deliver outstanding results on their own? To answer this question, in Tab. 10, we evaluate zero-shot retrieval to verify the importance of effectively utilizing narration information without training. For the experiment setting, we extracted CLIP features from each video and its narration, applied mean pooling, and evaluated zero-shot retrieval performance with given queries. In the table, the query-to-video retrieval shows reasonable retrieval performance without training, while the query-to-narration retrieval shows significantly lower performance. Although an element-wise matrix summation or the fusion approach used in our NarVid framework shows some improvements, the retrieval performances remain insufficient for practical retrieval scenarios. These experimental results once again support the claim of our framework that the information from the generative model should be utilized appropriately, including training.
@@
Temporal modeling Text-to-video retrieval
$\phi_{temp}(\boldsymbol{\hat{v}})$ $\phi_{temp}(\boldsymbol{\hat{n}})$ R@1 R@5 R@10
49.1 75.9 84.8
50.2 75.8 85.4
51.0 76.4 85.2
Table 12: Comparison of the impact of temporal modeling on performance across modalities on MSR-VTT
D.2 Temporal Modeling Details.
The temporal block in Eq. 3 is defined as $\phi_{temp}(\boldsymbol{\hat{v}})=Transformer(\boldsymbol{\hat{v}}+\boldsymbol{P})+\boldsymbol{\hat{v}}$, where $\boldsymbol{P}$ represents the positional embedding, as used in CLIP4Clip [25]. Temporal modeling is commonly applied to videos, as supported by Row 1 of Tab. 12. Similarly, although narrations have inherent temporal orders, we believe that effectively capturing the temporal relationships among frame-level captions enhances representation. Notably, removing $\phi_{temp}$ from the narration led to a 0.8% decrease in R@1, highlighting the significance of temporal modeling.
D.3 Cross-View Hard Negative Loss
Figure 8: Ablation study results of $\lambda$ and $\eta$, used in the hard negative loss, with $\alpha$ set to 1, conducted on the MSR-VTT.
Fig. 8 shows the performance of R@1 with varying values of $\lambda$ and $\eta$, which are hyperparameters used in the hard negative rank loss of Eqs. 6 and 8.
$\lambda$ is a scaling factor that establishes the threshold for selecting hard negatives within the query-video and query-narration similarity matrices. Likewise, $\eta$ serves as a scaling factor that modifies the thresholds in the cross-view hard negative loss. The balance between $\lambda$ and $\eta$ is crucial for optimal performance, achieving its best results at $\lambda=0.7$ and $\eta=1.8$.
Figure 9: Variations in selected hard negatives with the increase of $\lambda$, along with the corresponding recall results on MSR-VTT.
Figure 10: Variations in selected hard negatives with the increase of $\alpha$, along with the corresponding recall results on MSR-VTT.
Intuitively, employing the same hyperparameters during training may lead to a decrease in the selection of hard negative samples, consequently diminishing the impact of the hard negative loss. This raises an intriguing question about the benefits of focusing on hard negative samples in the later stages of training. To address this question, we linearly increased $\lambda$ to acquire more hard negative samples during experimentation. As shown in Fig. 9, our findings reveal that while this approach successfully increased the quantity of hard negative samples, it resulted in a degradation of performance. This indicates that forcing easy samples to be treated as hard negatives may hinder representation learning.
Fig. 10 presents the results of linearly increasing $\alpha$, which similarly shows performance degradation. Notably, in the case of increasing $\alpha$ from 1.0 to 2.0, which emphasizes the hard negative loss, we found that the number of selected hard negative samples actually increased. The result suggests that dynamically modifying the hard negative loss weight to focus more on hard negatives may negatively impact the learned representations.
@@ D.3 Cross-View Hard Negative Loss
$L_{qv}^{H}=-\frac{1}{2B}\sum_{i=1}^{B}\left\{\log\frac{e^{\boldsymbol{S}_{qv}(i,i)}}{\displaystyle\sum_{{j}\in{H_{i}}}e^{\boldsymbol{S}_{qv}(i,j)}}+\log\frac{e^{\boldsymbol{S}_{qv}(i,i)}}{\displaystyle\sum_{{j}\in{H_{i}^{T}}}e^{\boldsymbol{S}_{qv}(j,i)}}\right\},$
$L_{qn}^{H}=-\frac{1}{2B}\sum_{i=1}^{B}\left\{\log\frac{e^{\boldsymbol{S}_{qn}(i,i)}}{\displaystyle\sum_{{j}\in{H_{i}}}e^{\boldsymbol{S}_{qn}(i,j)}}+\log\frac{e^{\boldsymbol{S}_{qn}(i,i)}}{\displaystyle\sum_{{j}\in{H_{i}^{T}}}e^{\boldsymbol{S}_{qn}(j,i)}}\right\},$
@@
D.4 Cost per Each Module.
For the number of trainable parameters, the two Cross-Modal Interactions have $\phi_{co-attn}$ (3.7M) and $\phi_{temp}$ (12.7M), respectively. The Narration Matching has 0.3M, leading to a 12% increase over the baseline [25] (136.2M). For FLOPs, the Narration Matching shows a 30.5G increase due to the expanded caption encoding, reaching 1.5 times the baseline (54.4G). The others have little impact.